find the measure of the larger angle of two complementary angles such that one half the larger is 15 more than one-fourth the smaller angle.
Answer by checkley77(12569)

Let x be the smaller angle & y be the larger angle.
y/2=15+x/4
x+y=90
y=90-x
(90-x)/2=15+x/4
(90-x)/2-x/4=15
[4(90-x)-2x]/2*4=15
(360-4x-2x)/8=15
360-6x=8*15
360-6x=120
-6x=120-360
-6x=-240
x=-240/-6
x=40 degrees is the smaller angle
40+y=90
y=90-40
y=50 degrees is the latger angle.
Proof:
50/2=15+40/4
25=15+10
25=25